Status Report

ISS On-Orbit Status 8 Mar 2002

By SpaceRef Editor
March 8, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

The ground expressed appreciation for the crew’s support during yesterday’s SSRMS (space station remote manipulator system) troubleshooting. They were also congratulated on the execution of the S0 dry-run, which was completed successfully while obtaining additional data to help MCC-H and CSA troubleshoot the SSRMS brake bus problem. After repeating a brakes release test several times, the failure in the Wrist Roll Joint (Prime String) was confirmed. Three other anomalies, apparently due to software problems, are being investigated.

In the payloads area, EarthKAM operations continue. FE-1 Carl Walz completed a change from the 50 mm lens to the 180 mm tele-lens, an activity which had to be performed during orbital night (5:10 am EST).

Today’s EVARM (EVA radiation monitoring) badge reading by Walz accomplished (approximately) the required “EVA minus 28-day” background data take. Badges are read on days 28, 21, 14, 7, 2, and 1 prior to an EVA. At the start of the EVARM activity, Carl needed to free up the HRF (human research facility) power converter for the EVARM reader, which required taking EarthKAM temporarily off and letting it continue its run on battery power. After EVARM was complete, deactivated and stowed, Walz restored the original power configuration for EarthKAM.

FE-2 Dan Bursch performed the first plant sampling of the ADVASC (Advanced Astroculture), reverting to taking pictures with the digital DCS 760 camera instead of collecting samples of the newly grown mustard plant seedlings as originally planned. Before the sampling, POIC (Payload Operations Integration Center) powered down ADVASC and terminated water flow. Comprehensive analysis of ADVASC science/health & status data and images has suggested that the plants have been developing slower than expected. Investigators have decided to defer actually cutting and preserving plant tissue and instead take some additional pictures of several views inside the growth chamber. These images should give us a better idea of how much the plants have developed and will help determine when the actual plant tissue sample activities should be rescheduled. Views captured with the digital camera included the entire root tray, a close up of each quarter of the root tray, a close up of the largest plant in the chamber, and “anything else that looks like it could be living tissue”. After the sample activity, ADVASC was powered up again, including reconnection of the air line to the growth chamber.

CDR Yuri Onufrienko continued maintenance of RS ventilation systems, today in the SM, while Walz completed work on the gas-liquid heat exchangers in the FGB, cleaning removable grills.

On request by MCC-Moscow, Onufrienko also took an inventory of all LIV Betacam and DVCAM videocassettes on board, bringing the IMS up to date, which showed 54 such cassettes.

At MCC-Houston, troubleshooting continues on the VRS (vacuum resource system) leak problem. By determining and reporting the current configuration of the VES/VRS jumper, the crew assisted in reducing the effort required for the procedure to isolate the location/cause of the vacuum leak. With the jumper configuration established, the ground now can now conduct leak tests of selected parts of the VRS.

During the recent utilization by the crew of the IWIS (internal wireless instrumentation system) to take structural dynamics measurements during physical exercise by Dan Bursch on the TVIS treadmill, good data were obtained by the ground, both by way of files and video camera recordings of the TVIS and the runner’s legs. Specialists are now starting to perform their analysis, and they expect good results.

The two flight engineers removed and replaced a failed BCM (battery charger module) in the Joint Airlock (JAL). The Avionics rack in the JAL contains four BCMs. The job took about 1.5 hours.

Monthly maintenance of the CSA-CP (compound specific analyzer – combustion products) was performed by FE-2 Bursch. This included exchanging batteries and switching prime, backup and sampling pumps to verify correct CO, HCN and HCl readings.

Onufrienko and Bursch collected fungal (mildew) samples from selected locations in the SM, wearing anti-aerosol respirators and goggles. Bursch used the U.S. SSK (surface sampler kit). Samples were carefully wrapped, labeled and sealed for return to Earth for analysis. Pictures were also taken of the affected areas with the DSC for downlink to MCC-M. If indicated by review of the sample in-flight analysis, the crew will treat the areas with Fungistat disinfectant.

After today’s successful conclusion of the fifth EVA on the STS-109/Hubble mission and reboost of the HST by the Columbia, a ship-to-ship comm interchange between ISS and STS-109 is planned for tomorrow for about 3:30 am EST, using both Ku and S-band. ISS onboard setup will be as for any live audio/video PAO event with the ground. This event will go out live on NASA-TV and will be replayed later.

ISS reboosts #5 and #6 of the current UF-1 stage with the Progress 6P are planned for 3/12 (next Tuesday). Burn 1 will be at 7:06 pm EST, for a delta-V of 2.0 m/s, burn 2 at 7:52 pm, for 3.5 m/s. These reboosts are designed by Moscow to set up orbit phasing for the 7P launch on 3/21 as well as the 8A rendezvous altitude. Precise maneuver data will be jointly finalized on 3/11.

Today’s CEO (crew earth observations) target areas were Mekong River Delta (detailed mapping views of this major delta system are needed to assess the increasing impact of human activity in both the delta itself and further upstream. Crew was to particularly note water color and agricultural activity while dry weather persists), E. Mediterranean Dust and Smog (blocking ridge of high pressure holds over the eastern Med, a favorable area for aerosol accumulation. As ISS approached the Sea of Marmara from the NW, crew was to try for oblique or limb shots to either side of track to help detect and quantify aerosol buildup in this region), Tropical Cyclone Hary (Indian Ocean)(strong tropical cyclone Hary is forecast to brush the NE coast of Madagascar with 85-kt winds about the time of this pass. Of interest: near-nadir views of the eye of this strengthening storm system), Congo-Zimbabwe Biomass Burning (dry weather is starting in Zimbabwe. Of interest: oblique or limb views to the left of track looking for smoke plumes or old burn scars), and Angolan Biomass Burning (high pressure has over southern Africa will provide a good off shore airflow this pass. As crew approached the Namibian coast from the NW, it was to look obliquely to the left track for aerosol plumes moving out to sea).

U.S. and Russian Segment Status (as of 2:15 pm EST today):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (32 Amps mode). Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is in Standby.
  • BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, #2 in Purify mode.
  • SM Working Compartment: Pressure (mmHg) — 759, temperature (deg C) — 27.2, ppO2 (mmHg) — 153.8, ppCO2 (mmHg) — 2.9.
  • SM Transfer Compartment: Pressure (mmHg) — 762, temperature (deg C) — 20.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • FGB Cabin: Pressure (mmHg) — 756, temperature (deg C) — 20.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Node: Pressure (mmHg) — 750.89, temperature (deg C) — 22.8 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 752.90, temperature (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 752.59, temperature (deg C) — n/a; shell heater temp (deg C) — 22.6, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1: Shell heater temp (deg C) — 21.9.
  • PMA-2: Shell heater temp (deg C) — 18.3.

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]).

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Directed (“parked”) position at 125 degrees.
  • SM batteries: battery #3 is cycling; all other batteries (7) in “partial charge” mode.
  • FGB battery #5 is offline; all other batteries (5) are in “partial charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is back-up.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • S-band is operating nominally.
  • Ku-band is operating nominally.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) operating nominally.


  • SSRMS/Canadarm2 at PMA-3 clearance position, with Keep Alive power on both strings.
  • RWS (robotics workstations) are Off.

ISS Orbit (as of this morning, 7:32 am EST):

  • Mean altitude — 386.8 km
  • Apogee — 389.9 km
  • Perigee — 383.5 km
  • Period — 92.3 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0004773
  • Orbits per 24-hr. day — 15.60
  • Altitude decrease — 410 m (mean) in last 24 hours
  • Solar Beta Angle: -26.0 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov.’98) — 18833
  • Current Flight Attitude — LVLH (local vertical/local horizontal = “earth-fixed” [yaw: -10 deg, pitch: -7 deg., roll: 0 deg])roll: 0 deg])

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.